Air streams horizontally past a small airplane's wings such that the speed is 70.0 m/s over the top surface and 50.0 m/s past the bottom surface. If the plane has a mass of 600 kg and the TOTAL area of the two wings is 9.00 m2, what is the net vertical force (including the effects of gravity) on the airplane? A. 3930 N B. 4920 N C. 6440 N D. 7390 N E. 9560 N

use p(air)=1.0
Answer #1

Solution :-

Applying Bernoulli equation below and above the wing:

P1 + 1/2*ρV12 = P2 + 1/2*ρV22

(P1 - P2) = 1/2*ρ(V22 - V12)

(P1 - P2) = 1/2*1.2*(702 - 502)

(P1 - P2) = 1440 Pa

Lift force L = (P1 - P2)*Area = 1440*9 = 12960 N

Downward force (weight) = 600*9.81 = 5886

Net force = 12960- 5886 = 7074 N.

This force acts downward.